OpenStudy (anonymous):

Find the perimeter of a quadrilateral with vertices at G (1, 6), H (2.5, 4), I (1, 2), K (-0.5, 4). Round your answer to the nearest hundredth when necessary. (5 points) Select one: a. 5 units b. 12.5 units c. 7.5 units d. 10 units

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@Nurali

OpenStudy (anonymous):

find the distance between all the adjacent points of the quadrilateral and add them up to get the perimeter of this quadrilateral

OpenStudy (nurali):

The perimeter is just the sum of the lengths of the sides. The lengths of the two horizontal sides (GH and IK) are just the differences in x-coordinates. For the lengths of the other sides, we'll need the distance formula based on the Pythagorean Theorem: the distance is the square root of the sums of the squares of the horizontal and vertical distances. The length of GH is [2.5 - (1)] = 1.5 The length of HI is √[(1-2.5)^2 + (2-4)^2] = √[2.25+4] = 2.5 The length of IK is [1 - (-0.5)] = 1.5 The length of GK is √[[1 - (-0.5)]^2 + (6-4)^2] = √[1.22+4] = 2.29 The perimeter is the sum: 1.5+2.5+1.5+2=7.5 units

OpenStudy (anonymous):

GH=2.5 IK=2.5 IH=2.5 GK=2.5 Perimeter =10 @nurali